# Explainable Artificial Intelligence for Autonomous Driving: A Comprehensive Overview and Field Guide for Future Research Directions

### v4

@@ Abstract
Autonomous driving has achieved significant milestones in research and development over the last two decades. There is increasing interest in the field as the deployment of autonomous vehicles (AVs) promises safer and more ecologically friendly transportation systems. With the rapid progress in computationally powerful artificial intelligence (AI) techniques, AVs can sense their environment with high precision, make safe real-time decisions, and operate reliably without human intervention. However, intelligent decision-making in such vehicles is not generally understandable by humans in the current state of the art, and such deficiency hinders this technology from being socially acceptable. Hence, aside from making safe real-time decisions, AVs must also explain their AI-guided decision-making process in order to be regulatory compliant across many jurisdictions. Our study sheds comprehensive light on the development of explainable artificial intelligence (XAI) approaches for AVs. In particular, we make the following contributions. First, we provide a thorough overview of the state-of-the-art and emerging approaches for XAI-based autonomous driving. We then propose a conceptual framework that considers all the essential elements for explainable end-to-end autonomous driving. Finally, we present XAI-based prospective directions and emerging paradigms for future directions that hold promise for enhancing transparency, trustworthiness, and societal acceptance of AVs.
keywords:
Explainable artificial intelligence , autonomous driving , intelligent transportation systems , regulatory compliance , safety
[…]
@@ 1 Introduction
A survey of the American National Highway Traffic Safety Administration (NHTSA) reports that nearly 94% of road accidents are due to human errors [1]. These human-related mistakes are mainly classified as driver distraction, drunk or impaired driving, lack of attention, violation of the traffic rules, limited view of traffic conditions, and jay-walking pedestrians [2]. The lack of rule obedience, the increasing number of vehicles on roads, and improper road culture have, therefore, motivated officials, manufacturers, and legislators to make substantial improvements in transportation systems. There are growing research and development attempts to enhance safety and automation capability of AVs, prevent traffic accidents, and create a better road infrastructure. In this sense, the potential benefits of AVs are improved convenience, operational safety (especially for seniors and people with reduced mobility) [3], reduced $CO_{2}$ emissions [4], diminished transportation costs [5], improved reliability [6, 7], and reduced traffic density [8]. Intel’s report on the projected benefits of AVs estimates that deployment of this technology on roads will result in a reduction of 250 million hours of users’ commuting time per year and save more than half a million lives from 2035 to 2045, just in the USA [9]. While the potential impact
Figure 1: A canonical exemplar of explainable AI in autonomous driving: An autonomous vehicle provides a natural and intelligible explanation of its real-time decision to bystanders. Graphics adapted and modified from the source: [10].
and benefits of AVs in everyday life are promising, there is a major societal concern about functional safety of such vehicles. This issue, as a major drawback, originates mainly from reports of recent traffic accidents with the presence of AVs, primarily owing to their opaque and inappropriate decision-making [11, 12, 13, 14]. As AI approaches provide the foundation for real-time driving actions, there is an inherent need and expectation from consumers and regulators that AI-driven decisions of AVs should be explainable  [^1] We’ll use the terms explainable and interpretable interchangeably. (e.g., Figure 1) to build confidence in these vehicles. In a recent study, we have described the fundamental concepts and process steps associated with XAI-based autonomous driving [15]. In this paper, we extend the scope of that work by discussing the following research questions in depth:
@@
    What are promising future XAI directions toward trustworthy, responsible, and publicly acceptable AVs?
@@
The rest of the article is structured as follows. In Section 2, we provide background information and the factors triggering the need for the emergence of XAI in autonomous driving. We then describe the concept of explanations in autonomous driving by analyzing 1) cross-disciplinary perspectives necessitating explanations, 2) the role of various types of explanations for diverse explanation recipients, and 3) construction methodologies for explanations in Section 3. We present a brief overview of frontier AI approaches powering the real-time action decisions of AVs in Section 4. In Section 5, we provide a comprehensive survey of state-of-the-art investigations on explainable AI-based autonomous driving. Motivated by current limitations and trends delineated in these works, we present a general design framework for explainable autonomous driving and show key components of such a framework in Section 6. Finally, we describe potential challenges and a road map toward safety and explainability of next-generation AVs as future directions and sum up the article.
@@ 2.1 AVs at a glance
AVs are intelligent vehicles capable of sensing their environment and mapping such sensing data to real-time driving decisions using an AI-based driving system. To discern, identify, and distinguish the objects in their operational surroundings, these vehicles fuse information from a variety of sensors that help make real-time driving decisions [16, 17]. The history of contemporary AVs goes back to 1988, when ALVINN (Autonomous Land Vehicle In a Neural Network), the first neural network-powered self-driving vehicle taking camera images with a laser range finder, was able to produce control commands for the road-following task [18]. Current AVs deployed on road networks have different levels of automation based on their in-vehicle technologies and intelligent capabilities. SAE International (previously known as the Society of Automotive Engineers) has defined six levels of autonomous driving [19]: Level 0 - No automation (a human driver is responsible for all critical driving tasks); Level 1 - Driving assistance (a vehicle has automated driving support such as acceleration/braking or steering, but the driver is responsible for all other possible driving operations); Level 2 - Partial automation (Advanced Driving Assistance Systems (ADAS) operations such as steering and acceleration/braking are available in this level); Level 3 - Conditional automation (a vehicle has more advanced features such as object/obstacle detection and can carry out the majority of driving operations); Level 4 - High automation (a vehicle can fulfill all possible driving operations in a geofenced area); and Level 5 - Full automation (a vehicle can perform all driving operations in any likely scenario, and no human intervention is required).
In general, there are two main approaches to building autonomous driving systems in terms of their AI-based learning architecture: modular and end-to-end pipelines [13, 20]. The modular pipeline consists of four primary and interconnected modules categorized as perception, localization, planning, and control (Figure 2, a). Perception is the sensing of an operational environment defined as a combination of two tasks: road surface extraction and on-road object detection [21]. Information for the purpose of perception can be obtained via an integrated sensor suite, such as LIDAR, RADAR, visible spectrum cameras, and ultrasonic sensors [17, 22]. The process of localization enables an autonomous vehicle to accurately determine its position in the sensor model of the world [23, 24]. The most effective way to get the position of an autonomous car is to use satellite navigation-based systems. Among such systems, the Global Navigation Satellite System (GNSS) and its most popular type, Global Positioning System (GPS), is a universal sensor to determine the global location of a car [25]. Planning is a complex operation that integrates components of route planning (i.e., selection of a route in the road work from the initial point to the final destination) and behavior planning (such as interactions with other vehicles, and people that may be met on a trajectory). Finally, control of an autonomous vehicle is the appropriate execution of chosen actions. In modern AVs, control is typically carried out through the ADAS software.
@@
Overall, the modular pipeline leverages various sensor suites and algorithms for each module. While being comprised of standalone components makes the modular system more explainable and debuggable, such an architecture propagates errors to the next component, and thus, the overall pipeline error becomes cumulative [20, 26, 27].
In contrast to a modular pipeline, end-to-end autonomous driving has recently emerged as a paradigm
Figure 2: Modular vs. end-to-end autonomous driving. In the modular pipeline, the described four operations are carried out subsequently to produce control commands while end-to-end driving directly inputs raw sensor data and produces control commands.
@@
shift in the design and development of AVs. End-to-end autonomous driving takes the raw sensor data as visual input and yields a control command for the vehicle (Figure 2, b) [20, 28]. Particularly, recent breakthroughs in deep learning and computer vision algorithms and the availability of rich sensor devices have been the primary reasons for automotive researchers to leverage the end-to-end learning approach. The advantage of an end-to-end pipeline over its counterpart is that it directly produces driving actions by unifying perception, localization, planning, and control as one combined machine learning (ML) task. Furthermore, computational efficiency is improved via shared backbones in end-to-end learning, and in this way, potential information loss in intermediate layers is also avoided [20, 29].
2.2 Existing issues
AI approaches, which are currently predominated by deep learning algorithms, have brought considerable improvements to many essential components of autonomous driving technology, including advances in perception, object detection, and planning. As the AI-powered driving systems of vehicles advance, the number of AVs deployed to road networks has proliferated significantly in many developed European countries, the US, and Canada over the last decade [30]. However, the aforementioned road accidents involving such cars have caused public skepticism, and many studies have attempted to underscore the current limitations and issues with the design, development, and deployment of autonomous cars on roads. For example, Fleetwood [31] has investigated public health and ethical issues arising with the use of autonomous driving. Their study provides an in-depth analysis of the health issues, especially with the Trolley problem examples [32, 33] (hitting a pedestrian on an icy road or a parked car; driving and hitting five people or changing the direction of the steering wheel and hitting an individual, etc.). Another critical aspect of the work is a concern for the potential rights and liabilities of passengers sitting inside an autonomous car (for example, by using such a car, does a passenger agree to face potential risks; does a passenger have the responsibility and liability to protect other road users if an accident happens?). That study concludes with four directions - clear and cross-disciplinary discussions amongst stakeholders, including a driving system’s action planning choice of an algorithm; enhancing society’s knowledge on the issues and limitations of autonomous driving; confirming society’s knowledge on solutions of the current issues and proper use of AVs; and developing faithful, rational, and monitored standards for public health experts’ attention.
@@ 2.2 Existing issues
Some studies have directly focused on the concept of ethical crashing (i.e., if crashing is inevitable, how to crash?) and the Trolley problem mentioned above. For instance, the Moral Machine experiment [34], a well-known and hotly debated experiment, investigates a general community’s preferences on applied Trolley problems (inevitable accident scenarios with binary outcomes) and states that “these preferences can contribute to developing global, socially acceptable principles for machine ethics.” However, further discussion on this issue condemns this opinion and draws attention to the lack of safety principles [35], which force deeper consideration of such dilemmas [36]. Sohrabi et al. [37] have explored the outcomes of autonomous driving technology on public health in an urban area. They determined that this technology can affect public health in thirty-two directions: more than half of these pathways (seventeen) are negative, eight are positive, and the remaining ones are unsettled. Another similar study led by Martinho et al. [38] has inspected autonomous vehicle-related ethics in both scientific literature and various industry reports in California. According to this survey, both the scholarly literature and industry documents highlight safety and cybersecurity as main issues and underline concerns for moral decision-making algorithms, human-involved control, ethics, design, privacy, accountability, and sustainability, among other concerns. Burton et al. [39] have identified three open problems in the state-of-the-art development of autonomous systems. The first one is the semantic gap that emerges when a thorough specification of the system is not provided to manufacturers and designers. Another identified issue is the responsibility gap, which arises when an accident happens and the responsibility of either an autonomous system or a human is the cause of this accident remains unresolved. Finally, there is the question of who is responsible for compensating the injured during an accident, which precipitates the third issue: the liability gap. That study also shows that the core of these issues is associated with domain complexity, system complexity, and transferring more decision-making functions from humans to autonomous systems. In a relevant case study, authors relate unpredictable urban environments, driving system complexity, and lack of human override in highly autonomous driving to the enumerated core reasons, respectively. Finally, several studies have attempted to elucidate other related issues such as privacy [40, 41], design and implementation issues [42], legal regulation [43], user concerns [44], IoT challenges [45], integration in smart cities [46], as well as management [47] and security concerns [48]. The key findings outlined in the above studies require an understanding of the causes of these issues and intrinsically give the stakeholders the right to ask “why” questions. So, we immediately observe an immense need for explanations about the performance of self-driving cars. Providing explanations of critical decisions can significantly increase the acceptance of AVs by both the transportation jurisdictions and general community. In the subsequent sections, we discuss established standards and regulations for autonomous driving technology and describe the design principles of AI software architecture for it. We then analyze the need for explanations in AVs, emerging from these multidisciplinary points of view.
2.3 AVs regulations and standards
@@
The issues and growing concerns caused by AI systems create the need to scrutinize the regulation of this technology. As a result, public institutions have initiated the development of regulatory frameworks to monitor the activities of data-driven systems at both a country level and internationally. The focal points of these regulations are mainly to protect the stakeholders’ rights and ensure they have control over their data. For example, the General Data Protection Regulation (GDPR) of the European Union (EU) initiated guidelines to promote the “right of an explanation” principle for users, enacted in 2016 and taking effect in May 2018 [49]. Moreover, the EU has a specially defined strategy on Guidelines of Trustworthy AI that has seven essential requirements, namely 1) human agency, 2) technical robustness and safety, 3) privacy and data governance, 4) transparency, 5) accountability, 6) diversity, non-discrimination, and fairness, and 7) societal and environmental well-being; these principles are all to be applied in AI-based product research and development [50]. Similarly, Mohseni et al. [51] have tabled a broad description of ethical AI with a consequent impact on autonomous systems. In this context, AVs also must comply with these principles and provide intelligible explanations to the engaged stakeholders in support of an investigation of road accidents and other critical conditions involving AVs.
@@
Various organizations have recently proposed guidelines on the regulation of AVs to monitor their compliance with law enforcement. NACTO’s (National Association of City Transportation Officials) statement on automated vehicles proposes nine principles to shape a policy on regulation of future generation AVs [52]. Another well-defined guideline, The Research and Development (RAND) Corporation’s principles,
Figure 3: Multidisciplinary perspectives necessitating XAI for autonomous driving
covers promises and issues of AVs and an association of this technology to law and liability issues. Their principles also provide thorough guidance for public regulators to investigate transportation accidents and make safety recommendations (e.g., [53, 54]) to the regulators and manufacturers of AVs such as the American National Highway Traffic Safety Administration (NHTSA), SAE International, Tesla, and Apple. NHTSA of the US Department of Transportation has a specific federal guideline on automated vehicle policy to improve traffic safety [55] In March 2022, NHTSA announced that automobile manufacturers would no longer have to equip fully autonomous cars with manual control elements, such as steering wheel and braking pedals in the USA [56]. The Government of Canada has also recently released their comprehensive federal guidelines on testing and regulations of automated driving systems [57]. Their documentation provides detailed information and a regulatory road map for the relevant organizations on the engagement with government agencies, pre-trial, testing, and post-test considerations of AVs. In another recently adopted regulation, Germany has published an act on operations of driverless cars, particularly relevant to designated areas of public roads [58]. The UK government has also advanced their interests toward regulated and safe autonomous driving, and hands-free driving was expected to be legally allowed there by the end of 2021 [59]. Other developed countries such as Australia [60], and Japan [61] have also recently launched initiatives for trials of autonomous driving technology.
@@
While the regulations have been set out to ensure legislative norms and user demands are met, some standards provide specifications to achieve a high safety level, quality assurance, efficiency, and environmentally friendly transportation systems. The International Organization for Standardization (ISO) has adopted several standards to define the relevant issues on automated driving. Examples include the ISO 21448 [62], which specifies situational awareness standards to maintain operational safety under the “Safety of the Intended Functionality,” and the ISO 26262 [63] standard defined for the safety of electrical and electronic systems in production passenger vehicles, entitled as “Road vehicles – Functional safety.” In this context, ISO/TC 204 [64] is the primary standard that provides a comprehensive guide on the overall system and infrastructure aspects of intelligent transportation systems (ITS), supporting the standardization of AVs technology. Motivated by ISO/TC 204, some regional initiatives have also imposed relevant standards on the regulation of autonomous vehicles. For instance, The European Committee for Standardization, together with ISO, has the CEN/TC 278 [65] standard that develops acceptable levels of quality, use cases, and best practices for ITS in Europe. It turns out that AVs, or ITS, as a more general field, involve many multidisciplinary foundations to meet the involved stakeholders, insurance, and law enforcement requirements. Thorough documentation on the details of legislation, regulation, and standardization of automated vehicles can be viewed here [66].

### v5

@@ Abstract
Autonomous driving has achieved significant milestones in research and development over the last two decades. There is increasing interest in the field as the deployment of autonomous vehicles (AVs) promises safer and more ecologically friendly transportation systems. With the rapid progress in computationally powerful artificial intelligence (AI) techniques, AVs can sense their environment with high precision, make safe real-time decisions, and operate reliably without human intervention. However, intelligent decision-making in such vehicles is not generally understandable by humans in the current state of the art, and such deficiency hinders this technology from being socially acceptable. Hence, aside from making safe real-time decisions, AVs must also explain their AI-guided decision-making process in order to be regulatory compliant across many jurisdictions. Our study sheds comprehensive light on the development of explainable artificial intelligence (XAI) approaches for AVs. In particular, we make the following contributions. First, we provide a thorough overview of the state-of-the-art and emerging approaches for XAI-based autonomous driving. We then propose a conceptual framework that considers the essential elements for explainable end-to-end autonomous driving. Finally, we present XAI-based prospective directions and emerging paradigms for future directions that hold promise for enhancing transparency, trustworthiness, and societal acceptance of AVs.
Index Terms:
Explainable artificial intelligence, autonomous driving, intelligent transportation systems, safety, regulatory compliance.
@@ 1 Introduction
Asurvey of the American National Highway Traffic Safety Administration (NHTSA) reports that nearly 94% of road accidents are due to human errors [1]. Such a lack of rule obedience and improper road culture have, therefore, motivated officials, manufacturers, and legislators to make substantial improvements in transportation systems. In this sense, there are growing research and development attempts to enhance safety and automation capability of AVs with the goal of preventing traffic accidents, and creating a better road infrastructure. Intel’s report on the projected benefits of AVs estimates that deployment of this technology on roads will result in a reduction of 250 million hours of users’ commuting time per year and save more than half a million lives from 2035 to 2045, just in the USA [2].
While the potential impact and benefits of AVs in everyday life are promising, there is a major societal concern about functional safety of such vehicles. This issue, as a major drawback, originates mainly from reports of recent traffic accidents with the presence of AVs, primarily owing to their “black-box” decision-making [3, 4, 5, 6]. As AI approaches provide the foundation for real-time driving actions, there is an inherent need and expectation from consumers, general society, and regulatory bodies that AI-based action decisions of AVs should be explainable to build confidence in these vehicles [3, 7, 8, 9] (e.g., Figure 1).
In this survey, we present a comprehensive overview of state-of-the-art investigations on the explainability of autonomous driving.
Figure 1: A canonical example of explainable AI in autonomous driving: An autonomous vehicle provides a live natural language explanation of its real-time decision to bystanders. The image has been adapted and modified from the original source: [10].
Through extensive analyses, we first show the background information on the need for the emergence of explanations for AVs. Furthermore, we fill the gap in the current literature by providing a structured and comprehensive review of state-of-the-art and emerging XAI approaches for autonomous driving and present a road map for future directions. More specifically, we discuss the following research questions in depth:
@@
    What are promising future XAI directions toward trustworthy,
    Figure 2: Modular vs. end-to-end autonomous driving. In the modular pipeline, the described operations are carried out subsequently to produce control commands while end-to-end driving directly inputs raw sensor data and produces control commands as a unified task.
    responsible, regulatory compliant, and publicly acceptable AVs?
@@
The rest of the article consists of six sections. Section 2 provides background information and the factors triggering the need for the emergence of XAI in autonomous driving. Section 3 covers the concept of explainability for AVs by analyzing 1) cross-disciplinary perspectives necessitating explanations, 2) the role of various types of explanations for diverse explanation recipients, and 3) construction methodologies for explanations. Section 4 provides a comprehensive survey of studies on XAI-based autonomous driving. Motivated by current limitations and trends delineated in these works, Section 5 presents a general design framework for explainable autonomous driving and shows key components of such a framework. Finally, Section 6 outlines potential challenges and a road map toward safety and explainability of next-generation AVs, as future directions, and Section 7 concludes the paper with an overall summary.
@@ 2.1 Autonomous driving at a glance
AVs, also referred to as self-driving vehicles, are intelligent vehicles equipped with advanced sensors, cameras, RADAR, LIDAR, GPS, and sophisticated learning algorithms that enable them to navigate and operate without human intervention [11]. To discern, identify, and distinguish the objects in their operational surroundings, these vehicles fuse information from a variety of sensors that help make real-time driving decisions [12, 13]. The history of contemporary AVs goes back to 1988, when ALVINN (Autonomous Land Vehicle In a Neural Network), the first neural network-powered self-driving vehicle taking camera images with a laser range finder, was able to produce control commands for the road-following task [14]. Current AVs deployed on road networks have different levels of automation based on their in-vehicle technologies and intelligent capabilities. SAE International (previously known as the Society of Automotive Engineers) has defined six levels of autonomous driving [15]: Level 0 - No automation (a human driver is responsible for all critical driving tasks); Level 1 - Driving assistance (a vehicle has automated driving support such as acceleration/braking or steering, but the driver is responsible for all other possible driving operations); Level 2 - Partial automation (Advanced Driving Assistance Systems (ADAS) operations such as steering and acceleration/braking are available in this level); Level 3 - Conditional automation (a vehicle has more advanced features such as object/obstacle detection and can carry out the majority of driving operations); Level 4 - High automation (a vehicle can fulfill all possible driving operations in a geofenced area); and Level 5 - Full automation (a vehicle can perform all driving operations in any likely scenario, and no human intervention is required).
@@
There are two main approaches to building autonomous driving systems in terms of their AI-based learning architecture: modular and end-to-end pipelines [5, 16]. The modular pipeline consists of four primary and interconnected modules categorized as perception, localization, planning, and control (Figure 2, a). The modular pipeline leverages various sensor suites and algorithms for each module. While being comprised of standalone components makes the modular system more explainable and debuggable, such an architecture propagates errors to the next component, and thus, the overall pipeline error becomes cumulative [16, 17, 18].
@@
In contrast to a modular pipeline, end-to-end autonomous driving has recently emerged as a paradigm shift in the design and development of AVs. End-to-end autonomous driving takes the raw sensor data as visual input and yields a control command for the vehicle (Figure 2, b) [16, 19, 20]. Particularly, recent breakthroughs in deep learning and computer vision algorithms, and the availability of rich sensor devices along with enhanced safety benefits have been the primary reasons for automotive researchers to leverage the end-to-end learning approach. The advantage of an end-to-end pipeline over its counterpart is that it directly produces driving actions by unifying perception, localization, planning, and control as one combined machine learning (ML) task. Furthermore, computational efficiency is improved via shared backbones in end-to-end learning, and in this way, potential information loss in intermediate layers is also avoided [16, 21].
2.2 Fundamental issues
@@
AI approaches, which are currently predominated by deep learning algorithms, have brought considerable improvements to many essential components of autonomous driving technology, including advances in perception, object detection, and planning. As the AI-powered driving systems of vehicles advance, the number of AVs deployed to road networks has proliferated significantly in many developed European countries, the US, and Canada over the last decade [22]. However, the aforementioned road accidents involving such cars have caused public skepticism, and many studies have attempted to underscore the current limitations and issues with the design, development, and deployment of AVs on roads. For example, Fleetwood [23] has investigated public health and ethical issues arising with the use of autonomous driving. Their study provides an in-depth analysis of the health issues, especially with the Trolley problem examples [24, 25] (hitting a pedestrian on an icy road or a parked car; driving and hitting five people or changing the direction of the steering wheel and hitting an individual, etc.). Some studies have directly focused on the concept of ethical crashing (i.e., if crashing is inevitable, how to crash?) and the Trolley problem mentioned above. For instance, the Moral Machine experiment [26], a well-known and hotly debated experiment, investigates a general community’s preferences on applied Trolley problems (inevitable accident scenarios with binary outcomes) and states that “these preferences can contribute to developing global, socially acceptable principles for machine ethics.” However, further discussion on this issue condemns this opinion and draws attention to the lack of safety principles [27], which force deeper consideration of such dilemmas [28]. Burton et al. [29] have identified three open problems in the state-of-the-art development of autonomous systems. The first one is the semantic gap that emerges when a thorough specification of the system is not provided to manufacturers and designers. Another identified issue is the responsibility gap, which arises when an accident happens and the responsibility of either an autonomous system or a human is the cause of this accident remains unresolved. Finally, there is the question of who is responsible for compensating the injured during an accident, which precipitates the third issue: the liability gap. That study also shows that the core of these issues is associated with domain complexity, system complexity, and transferring more decision-making functions from humans to autonomous systems. Further studies include the outcomes of autonomous driving technology on public health in an urban area [30], and ethical dilemmas with AVs [31]. Overall, the key findings from these studies necessitate an understanding of the causes of these issues and intrinsically give the stakeholders the right to ask “why” questions.
2.3 Regulations and standards
@@
The issues and growing concerns caused by AI systems create the need to scrutinize the regulation of this technology. As a result, public institutions have initiated the development of regulatory frameworks to monitor the activities of data-driven systems at both a country level and internationally. The focal points of these regulations are mainly to protect the stakeholders’ rights and ensure they have control over their data. For example, the General Data Protection Regulation (GDPR) of the European Union (EU) initiated guidelines to promote the “right of an explanation” principle for users, enacted in 2016 and taking effect in May 2018 [32]. Moreover, the EU has a specially defined strategy on Guidelines of Trustworthy AI that has seven essential requirements, namely 1) human agency, 2) technical robustness and safety, 3) privacy and data governance, 4) transparency, 5) accountability, 6) diversity, non-discrimination, and fairness, and 7) societal and environmental well-being; these principles are all to be applied in AI-based product research and development [33].
@@
Various organizations have recently proposed guidelines on the regulation of AVs to monitor their compliance with law enforcement. NACTO’s (National Association of City Transportation Officials) statement on automated vehicles proposes nine principles to shape a policy on regulation of future generation AVs [34]. NHTSA of the US Department of Transportation has a specific federal guideline on automated vehicle policy to improve traffic safety [35]. In March 2022, NHTSA announced that automobile manufacturers would no longer have to equip fully autonomous cars with manual control elements, such as a steering wheel and braking pedals in the USA [36]. Canada [37], Germany [38], UK [39], Australia [40], and Japan [41] have also recently launched their regulations on autonomous driving technology.
@@
While the regulations have been set out to ensure legislative norms and user demands are met, some standards provide specifications to achieve a high safety level, quality assurance, efficiency, and environmentally friendly transportation systems. The International Organization for Standardization (ISO) has adopted several standards to define the relevant issues on automated driving. Examples include the ISO 21448 [42], which specifies situation awareness standards to maintain operational safety under the “Safety of the Intended Functionality,” and the ISO 26262 [43] standard defined for the safety of electrical and electronic systems in production passenger vehicles, entitled as “Road vehicles – Functional safety.” Thorough documentation on the details of legislation, regulation, and standardization of autonomous cars can be viewed in [44].